OpenStudy (anonymous):

A company did a quality check on all the packs of nuts it manufactured. Each pack of nuts is targeted to weigh 18.25 oz. A pack must weigh within 0.36 oz of the target weight to be accepted. What is the range of rejected masses, x, for the manufactured nuts? x < 17.89 or x > 18.61 because |x - 18.25| > 0.36 x < 17.89 or x > 18.61 because |x - 0.36| + 18.25 > 0 x < 18.25 or x > 18.61 because |x - 18.25| > 0.36 x < 18.25 or x > 18.61 because |x - 0.36| + 18.25 > 0

OpenStudy (anonymous):

Solve: -2(2x + 5) - 3 = -3(x - 1) 16 -16 -10 -13

OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 2) = 4x. negative 2 over 5 negative 5 over 2 2 over 5 5

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

That last one...is it... \[\large\frac{2}{3}(x - 2) = 4x\] ?

OpenStudy (anonymous):

Yes It Is Sorry^.^

OpenStudy (anonymous):

The table shows the solution to the equation |2x - 3| - 1 = 2: Step 1 |2x - 3| = 2 + 1 Step 2 |2x - 3| = 3 Step 3 2x - 3 = 3 or 2x + 3 = 3 Step 4 2x = 6 or 2x = 0 Step 5 x = 3 or x = 0 Which is the first incorrect step? Step 1 Step 3 Step 5 Solution is correct